### v1

# RGB-Only Supervised Camera Parameter Optimization in Dynamic Scenes

Abstract
Although COLMAP has long remained the predominant method for camera parameter optimization in static scenes, it is constrained by its lengthy runtime and reliance on ground truth (GT) motion masks for application to dynamic scenes. Many efforts attempted to improve it by incorporating more priors as supervision such as GT focal length, motion masks, 3D point clouds, camera poses, and metric depth, which, however, are typically unavailable in casually captured RGB videos. In this paper, we propose a novel method for more accurate and efficient camera parameter optimization in dynamic scenes solely supervised by a single RGB video. Our method consists of three key components: (1) Patch-wise Tracking Filters, to establish robust and maximally sparse hinge-like relations across the RGB video. (2) Outlier-aware Joint Optimization, for efficient camera parameter optimization by adaptive down-weighting of moving outliers, without reliance on motion priors. (3) A Two-stage Optimization Strategy, to enhance stability and optimization speed by a trade-off between the Softplus limits and convex minima in losses. We visually and numerically evaluate our camera estimates. To further validate accuracy, we feed the camera estimates into a 4D reconstruction method and assess the resulting 3D scenes, and rendered 2D RGB and depth maps. We perform experiments on 4 real-world datasets (NeRF-DS, DAVIS, iPhone, and TUM-dynamics) and 1 synthetic dataset (MPI-Sintel), demonstrating that our method estimates camera parameters more efficiently and accurately with a single RGB video as the only supervision.
Figure 1: (a) Overview of our RGB-only supervised camera parameter optimization. (b) Front view of the 3D Gaussian field reconstructed by our camera estimates at time $t$. (c) 2D renderings (RGB and depth) at time $t$ with quantitative metrics. Our optimization is not only significantly more efficient and accurate, but also avoids overfitting the reconstruction to specific viewpoints. Record3D is a mobile app that factory-calibrates the intrinsic and uses LiDAR sensors to collect metric depth for camera pose estimates, thus does not have valid runtime.
1 Introduction
Despite recent progress in visual odometry, efficiently and accurately optimizing camera parameters [^1] Like all existing methods in table 1, we also assume a pinhole camera. (focal length + rotation&translation) from casually collected RGB dynamic-scene videos remains a big challenge. Although the most predominant COLMAP colmap method [^2] We denote the COLMAP using motion masks as $\texttt{COL}^{\text{w/ mask}}$ and the one w/o motion masks as $\texttt{COL}^{\text{w/o mask}}$. is RGB-only supervised, it suffers from its lengthy runtime and requisite of GT motion masks to mask out the outlier moving stuff. In Table˜1, most recent approaches cf3dgs ; gflow ; leapvo ; vggsfm ; particlesfm ; dust3r ; monst3r ; cut3r attempted to improve through being supervised by additional GT priors such as focal length, metric depth, 3D point clouds, camera poses, and motion masks, which are typically unavailable in casually collected videos. We cannot help but ask a natural question: Is it possible to accurately and efficiently estimate camera parameters in dynamic scenes in an RGB-only supervised manner - the most minimal form of supervision?
Existing RGB-only supervised methods vggsfm ; gflow ; rodynrf ; particlesfm ; leapvo make obvious improvements, but they mostly rely on multiple pre-trained dense prediction models raft ; cotracker ; midas to compensate for the inaccuracies of individual pseudo-supervision sources, resulting in performance degradation if any of them fails. They also cannot adaptively exclude moving outliers without GT motion supervision. Besides, their high computational latency always leads to lengthy runtimes. Further discussion of related work is provided in Section˜2.
Table 1: Categorization of supervision of current methods. Ours, casualSAM casualsam , and Robust-CVD robustcvd are RGB-only supervised, while our performance is the best as shown in section˜4.
Supervision
Static Scene
Dynamic Scene
GT 3D Point Cloud & Camera Pose
Dust3r dust3r , Fast3r fast3r , Mast3r mast3r , Spann3r spann3r , VGGT vggt Monst3r monst3r , Cut3r cut3r , Stereo4D stereo4d , Easi3r easi3r
GT Focal Length
CF-3DGS cf3dgs , Nope-NeRF nope-nerf , LocalNeRF localnerf
    + Metric Depth
DROID-SLAM droidslam
      + GT Motion Priors
GFlow gflow , LEAP-VO leapvo
GT Motion Priors
RoDynRF rodynrf , $\texttt{COL}^{\text{w/ mask}}$ colmap , ParticleSfM particlesfm
RGB-Only
[…]
Based on these insights, we propose an RGB-only supervised, accurate, and efficient camera parameter optimization method, with a brief performance overview in Figure˜1. Specifically, to minimize reliance on pre-trained dense prediction models while still establishing robust and maximally sparse hinge-like relations across the video as accurate pseudo-supervision (bottom right corner in Figure˜2), we propose the novel patch-wise tracking filters built solely on a pre-trained point tracking (PT) model. This formulation effectively avoids inaccurate tracking trajectories extracted across frames and computational latency induced by the noisy dense prediction as pseudo-supervision.
However, the extracted pseudo-supervision includes a portion of trajectories belonging to moving outliers. To eliminate the influence of such outliers, we introduce a learnable uncertainty associated with each calibration point, where each is a learnable 3D position in the world coordinates, corresponding to one extracted tracking trajectory. We model such uncertainty parameters with the Cauchy distribution, which can deal with heavy tails better than, e.g., the Gaussian distribution, and propose the novel Average Cumulative Projection error and Cauchy loss for the outlier-aware joint optimization of the calibration points, focal length, rotation, translation, and uncertainty parameters. Unlike casualSAM and LEAP-VO, which assign uncertainty parameters to 2D pixels, our approach associates uncertainties with sparse 3D calibration points, resulting in significantly fewer learnable parameters and reduced runtime, as shown in LABEL:tab:average_time.
Such joint optimization is prone to getting trapped in local minima. To address this, we analyze the asymptotic behavior of the Softplus function and the analytical minima of the inner convex term in losses to propose a two-stage optimization strategy to accelerate and stabilize the optimization. We evaluate the performance of our method through extensive experiments on 5 popular public datasets - NeRF-DS nerfds , DAVIS davis , iPhone iphone , MPI-Sintel mpi-sintel , and TUM-dynamics tum-dynamics , demonstrating our superior performance. Our contributions can be summarized as follows.
  • We propose the first RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes with three key components: (1) patch-wise tracking filters; (2) outlier-aware joint optimization; and (3) a two-stage optimization strategy.
  • We present exhaustive quantitative and qualitative experiments and extensive ablation studies that demonstrate the superior performance of our proposed method and the contribution of each component.
2 Related Works
Dynamic Scene Reconstruction/Novel View Synthesis (NVS). Existing methods for reconstructing objects and scenes use a variety of 3D representations, including planar planer1 ; planer2 , mesh banmo ; limr , point cloud pointnerf ; pt2 , neural field nerf ; nerfds ; nerfstudio ; dnerf ; nerfinthewild , and the recently introduced Gaussian explicit representations gsslam ; 3dgs ; 2dgs ; 4dgs ; d3dgs . NeRF nerf enables high-fidelity NVS. Some methods nerfies ; hypernerf ; dnerf ; nerfds ; infonerf ; regnerf also extend NeRF to dynamic scenes, while others banmo ; limr ; volsdf ; neus ; nerfstudio build on them to extract high-quality meshes. However, NeRF-based methods have the limitation of a long training time. Recently, 3DGS 3dgs effectively addressed this issue by using 3D Gaussian-based representations and presented Differential-Gaussian-Rasterization in CUDA. 3DGS optimizes 3D Gaussian ellipsoids as dynamic scene representations associated with attributes such as position, orientation, opacity, scale, and color. Several studies 4dgs ; d3dgs also have used 3DGS for dynamic scenes, achieving near real-time dynamic scene novel view synthesis. However, both NeRF-based and 3DGS-based methods heavily rely on $\texttt{COL}^{\text{w/ mask}}$ to estimate camera parameters.
Camera Parameter Optimization. Many efforts have been made to overcome the shortcomings of COLMAP, particularly for dynamic scenes. But each suffers from some constraints. In Table˜1, we present a categorization of supervision of current SOTA methods. Supervised by additional GT focal length, CF-3DGS cf3dgs , Nope-NeRF nope-nerf , and LocalNeRF localnerf leverage a pre-trained monocular depth estimation model midas to estimate camera poses and the static scene jointly. The most representative SLAM-based method - DROID-SLAM droidslam , leverages both GT focal length and metric depth as supervision. GFlow and LEAP-VO gflow ; leapvo extend it to dynamic scenes with both GT focal length and motion priors as supervision. Although VGGSfM, FlowMap, InstantSplat, $\texttt{COL}^{\text{w/o mask}}$ vggsfm ; flowmap ; instantsplat ; colmap eliminate the GT focal length requirement by leveraging pre-trained PT models raft ; cotracker , they cannot handle the moving objects in dynamic scenes. RoDynRF rodynrf , $\texttt{COL}^{\text{w/ mask}}$ colmap , and ParticleSfM particlesfm simply tackle such a problem by incorporating GT motion supervision like GFlow. Recently, DUSt3R-based methods dust3r ; fast3r ; mast3r ; spann3r ; vggt and their dynamic-scene counterparts monst3r ; cut3r ; stereo4d ; easi3r explored feed-forward camera parameter prediction by training on large-scale static and dynamic scene datasets, respectively, in a fully supervised manner - that is, using GT 3D point clouds and camera poses as supervision, requiring several days’ training on high-end GPUs. However, unlike LLMs that benefit from abundant language data, such metric 3D supervision is relatively scarce in the vision area, leading to frequent domain gaps when these models are applied to unseen data. In contrast, Robust-CVD robustcvd , casualSAM casualsam and our method conduct camera parameter optimization for dynamic scenes in a more general RGB-only supervised way. However, as shown in Section˜4, their performance is significantly worse than ours.
3 Method
Under RGB-only supervision, RGB frames $F_{i},i\in[0,N-1]$ (NN is frame count) are given. Our proposed patch-wise tracking filters (Section˜3.1) extract $H$ robust and maximally sparse hinge-like tracking trajectories as pseudo-supervision, where each corresponds to one calibration point $\mathbf{P}_{h}^{cali}\in\mathbb{R}^{3},h\in[0,H]$ in the world coordinates. Under such pseudo-supervision and our newly proposed ACP error and Cauchy loss, the calibration points $\mathbf{P}^{cali}$, focal length $f\in\mathbb{R}$, quaternion matrix $\mathbf{Q}\in\mathbb{R}^{N\times 4}$, translation $\mathbf{t}\in\mathbb{R}^{N\times 3}$, and motion-caused uncertainty parameters $\mathbf{\Gamma}\in\mathbb{R}_{>0}^{H}$ are jointly optimized (Section˜3.2). $\mathbf{\Gamma}$ is the scale parameter of the Cauchy distribution which is used to model such uncertainty parameters, associated with each calibration point, to reduce the erroneous influence of moving outliers. By analyzing the Softplus limits and convex minima in losses, we propose a simple but effective two-stage optimization strategy (Section˜3.3) to enhance the stability and optimization speed.
3.1 Patch-wise Tracking Filters
Built on a pre-trained PT model, we observe that its attention mechanism assigns higher attention weights to pixels with more accurate tracking results which are always texture-rich pixels with large gradient norms. Inspired by it, as shown in Figure˜2, we propose the patch-wise texture filter to identify the high-texture patches within $F_{t}$ and the patch-wise gradient filter to select the pixel with the highest gradient norm within each identified patch. While tracking such identified points, the visibility filter keeps removing trajectories that become invisible and the patch-wise distribution filter keeps the one with the largest gradient norm when multiple moving points enter the same patch. As shown in section˜E.2.1 (fig.˜9), our method only retains the robust and accurate trajectories as pseudo-supervision.
Figure 2: Patch-wise tracking filters. (1) Partitioning $F_{0}$ into patches of size $w\times w$, the patch-wise texture filter computes the texture map $\mathbf{T}_{0}$ and marks the high-texture patches in gray; (2) Within each high-texture patch, the patch-wise gradient filter selects one potential tracking point with the highest gradient norm. (3) The visibility filter removes the entire trajectory of a point if it becomes invisible at any time ($\bullet$, $\blacksquare$, $\blacktriangle$ $\rightarrow$ kept trajectories; $\circ$, $\square$, $\triangle$ $\rightarrow$ removed trajectories); (4) The patch-wise distribution filter only keeps the one with the largest gradient norm when multiple trajectories fall into the same patch. $\mathbf{P}$ and $\mathbf{I}$ are the location and index of trajectory, and $\leftrightarrow$ is the trajectory range.
Patch-wise Texture Filter. Highly distinguishable points, that can be tracked reliably, belong to highly nonuniform (textured) neighborhoods. To identify such neighborhoods, our patch-wise texture filter computes a texture map $\mathbf{T}_{i}\in\mathds{1}^{\mathbb{H}/w\times\mathbb{W}/w}$, giving a measure of texture level for each $w\times w$ patch where $\mathbb{H}$ and $\mathbb{W}$ denote the height & width of $F_{i}$. We represent the texture level of a patch by
$\mathbf{T}_{i}[m,n]=\mathds{1}\{\Sigma_{i}[m,n]>\tau_{var}\cdot\sigma^{*}\}$ (1)
, where $\Sigma_{i}\in\mathbb{R}^{\mathbb{H}/w\times\mathbb{W}/w}$ is the intensity variance, $\sigma^{*}=\max(\Sigma_{i})$, $\tau_{var}$ is the percentage threshold of minimum variance for the patch to be selected, and $m,n\in[0,\mathbb{H}/w-1],[0,\mathbb{W}/w-1]$. The texture levels of a patch are represented by 1 for the selected patches and 0 for the others.
Patch-wise Gradient Filter. Within the identified patches, our patch-wise gradient filter computes the intensity gradient norm map $\mathbf{G}_{i}\in\mathbb{R}^{\mathbb{H}\times\mathbb{W}}$ of $F_{i}$, and selects the point with the largest gradient norm within each patch. This yields the pool of potentially distinguishable points, forming potential trajectories, namely,
[…]
Visibility Filter. We find that current PT models cotracker ; cotracker3 ; sam-pt still tend to suffer from reduced tracking accuracy when a point becomes occluded and later reappears, due to the disruption of temporal feature continuity. Thus, if any $\tilde{\mathbf{P}}$ in any $F_{i}$ becomes invisible, our visibility filter deletes it by the dot product $\tilde{\mathbf{P}}\cdot\mathbf{V}$, $\mathbf{V}\in\{0,1\}\sim\tilde{\mathbf{P}}$, where $\mathbf{V}=0$ if a point is invisible.
Patch-wise Distribution Filter. This filter enforces a more even point distribution within each frame, preventing them from clustering into a small region as the viewpoint changes. It also helps reduce susceptibility to loss of resolution which might result in triangulation errors. We keep the highest-gradient tracking point $\tilde{\mathbf{P}}^{*}$ in each patch $Pat_{m,n}$ of $F_{i}$, as follows:
$\tilde{\mathbf{P}}^{*}=\arg\max_{\tilde{\mathbf{P}}}\mathbf{G}_{i}[\tilde{\mathbf{P}}\in Pat_{m,n}],\;\text{if}\sum\mathds{1}(\tilde{\mathbf{P}}\in Pat_{m,n})>1$ (3)
As shown in Figure˜2, locations and indices of $\tilde{\mathbf{P}}^{*}$ are stored in $\mathbf{P}_{i}\in\mathbb{R}^{B\times 2}$ and $\mathbf{I}_{i}\in\mathbb{R}^{B}$, $i\in[0,H-1]$, acting as pseudo-supervision in the outlier-aware joint optimization. Each iteration starts at $F_{t},t=\arg\min_{t}(-1\in\mathbf{I}_{t})$, and ends until each frame contains exactly $B$ tracked points.
3.2 Outlier-aware Joint Optimization
Outlier-aware Joint Optimization Mechanism. Under the obtained pseudo-supervision, $\mathbf{P}_{cali}$, $f$, $\mathbf{Q}$, $\mathbf{t}$ and $\mathbf{\Gamma}$ are jointly optimized. We first project $\mathbf{P}^{cali-homo}\in\mathbb{R}^{H\times 4}$ (the homogeneous coordinates of $\mathbf{P}^{cali}$ obtained by concatenating $1$) onto each frame by
$\mathbf{P}_{i}^{proj-homo}=\mathbf{P}^{cali-homo}[\mathbf{I}_{i}]\cdot\begin{bmatrix}\mathbf{R}_{i}&\mathbf{t}_{i}\\ \mathbf{0}&1\end{bmatrix}^{T}\cdot\mathbf{K}^{T}\\$ (4)
$\mathbf{P}_{i}^{proj}=\mathbf{P}_{i}^{proj-homo}[:,:2]/\mathbf{P}_{i}^{proj-homo}[:,3]\\$ (5)
, where $i\in[0,N-1]$ and $\mathbf{P}^{proj}\in\mathbb{R}^{N\times B\times 2}$. $\mathbf{P}^{proj-homo}\in\mathbb{R}^{N\times B\times 4}$ denotes the homogeneous 2D location of the projection $\mathbf{P}^{proj}$. The perspective projection matrix $\mathbf{K}\in\mathbb{R}^{4\times 4}$ is derived from $f$, and the world-to-camera transformation matrix consists of rotation $\mathbf{R}_{i}$ and translation $\mathbf{t}_{i}$. We assume constant $f$ like SOTA particlesfm ; casualsam ; rodynrf . Notably, we learn the quaternion matrix $\mathbf{Q}_{i}$ instead of optimizing the $\mathbf{R}_{i}$ and additional constraints. This optimization approach circumvents the difficult-to-enforce orthogonality and $\pm 1$ determinant constraints required for rotation matrices during optimization.
Figure 3: Outlier-aware Joint Optimization. $\bullet$ represents $\mathbf{P}_{t}$ and $\mathbf{P}_{t^{\prime}}$ on each frame. The static samples $p^{\prime cali}$ and $p^{\prime\prime cali}$ can establish concrete triangulation relations with their corresponding $\mathbf{P}_{t}$, $\mathbf{P}_{t^{\prime}}$, and cameras, resulting in lower $\gamma^{\prime}$ and $\gamma^{\prime\prime}$. In contrast, the dynamic sample $p^{\prime\prime\prime cali}$ exhibits the opposite behavior.
Figure 4: Runtime Trends. As the frame count increases, our runtime grows almost linearly, whereas $\texttt{COL}^{w/omask}$ scales exponentially. The runtime of casualSAM is too large to fit in this figure. The complete runtime is in LABEL:tab:average_time, and section˜E.1.1 (table˜9, table˜10, table˜11).
However, the extracted pseudo-supervision always contains moving outliers. To mitigate its impact, without any GT motion priors, we identify such outliers by modeling the uncertainty their presence may cause in the observed distributions of the inlier points. We introduce the uncertainty $\mathbf{\Gamma}\in\mathbb{R}^{H}$ associated with $\mathbf{P}^{cali}\in\mathbb{R}^{H}$ and incorporate the Cauchy distribution $f(x;x_{0},\mathbf{\Gamma})=\frac{1}{\pi\mathbf{\Gamma}[1+(\frac{x-x_{0}}{\mathbf{\Gamma}})^{2}]},\quad\mathbf{\Gamma}>0$ to model the uncertainty parameter $\mathbf{\Gamma}$ since this distribution can better handle the heavy tails than, e.g., the Gaussian distribution. As depicted in Figure˜4, during optimization, inliers are expected to have low uncertainty, while outliers have high uncertainty. Since the scale parameter $\mathbf{\Gamma}$ in $f(x;x_{0},\mathbf{\Gamma})$ is required to be strictly positive, we introduce a new parameter $\mathbf{\Gamma}^{raw}$ which we obtain $\mathbf{\Gamma}$ from using the Softplus function $\mathbf{\Gamma}=\log(1+e^{\mathbf{\Gamma}^{raw}}),\mathbf{\Gamma}^{raw}\in\mathbb{R}^{H}$. This effectively ensures $\mathbf{\Gamma}\in\mathbb{R}_{>0}^{H}$ is differentiable and has smooth gradients.
Losses. To down-weight outliers by learned $\mathbf{\Gamma}$, we replace the commonly used projection error $\mathbb{E}^{proj}=\|\mathbf{P}^{proj}-\mathbf{P}\|_{2}^{2}$ with our proposed Average Cumulative Projection (ACP) error, defined as:
$\mathbb{E}_{h\in[0,H-1]}^{ACP}=\frac{\sum\mathds{1}_{\{\mathbf{I}=h\}}\circ\|\mathbf{P}^{proj}-\mathbf{P}\|_{2}^{2}}{\sum\mathds{1}_{\{\mathbf{I}=h\}}}$ (6)
, where $\mathbb{E}^{ACP}\in\mathbb{R}^{H}$ and $\circ$ denotes the element-wise matrix multiplication. For each $\mathbf{P}_{h}^{cali}$, we accumulate the errors between its corresponding projection and tracking locations across the video, then take the average as $\mathbb{E}_{h\in[0,H-1]}^{ACP}$. Furthermore, we propose the novel Cauchy loss $\mathcal{L}_{cauchy}$ in terms of the negative-log-likelihood $\log{(\mathbf{\Gamma}+\frac{(x-x_{0})^{2}}{\mathbf{\Gamma}})}$ of $f(x;x_{0},\mathbf{\Gamma})$ where we replace $x-x_{0}$ with $\mathbb{E}^{ACP}$ as eq.˜7. Our total loss $\mathcal{L}_{total}$ in Equation˜8 consists of $\mathcal{L}_{cauthy}$ and a depth regularization term $\mathcal{R}_{depth}$ to encourage positive depth. With the estimated camera parameters, we use 4DGS 4dgs for scene reconstruction. Reconstruction and loss derivation details are in appendix˜A and appendix˜B.
$\mathcal{L}_{cauchy}=\frac{1}{H}\sum_{h=0}^{H}\log{(\mathbf{\Gamma}+\frac{(\mathbb{E}^{ACP})^{2}}{\mathbf{\Gamma}})}$ (7)
$\mathcal{L}_{total}=\mathcal{L}_{cauthy}+\mathcal{R}_{depth},\;\mathcal{R}_{depth}=\frac{1}{N}\sum_{i=0}^{N}-\text{ReLU}(\mathbf{P}_{i}^{proj-homo}[:,3])$ (8)
3.3 Two-stage Optimization Strategy
To avoid convergence to local minima, we propose this strategy based on an analysis of the asymptotic behavior of the Softplus function and the analytical minima of the inner convex term in $\mathcal{L}_{cauthy}$. Stage 1 focuses on rapid convergence, while Stage 2 aims for stable convergence by initializing $\mathbf{\Gamma}^{raw}$ to the ACP error after Stage 1. The effectiveness of it is concretely demonstrated in Figure˜8.
Stage 1. In the Softplus function, $\mathbf{\Gamma}=\log(1+e^{\mathbf{\Gamma}^{raw}})\approx\mathbf{\Gamma}^{raw}$, as $\mathbf{\Gamma}^{raw}\to+\infty$. So in Stage 1, we fix $\mathbf{\Gamma}^{raw}=1$ and optimize only $\mathbf{P}^{cali}$, $f$, $\mathbf{Q}$, and $\mathbf{t}$ for quick convergence. The loss will converge to a certain value beyond the global minimum, as there is no proper $\mathbf{\Gamma}$ to down-weight outliers.
Stage 2. The inner term $\Phi=x+\frac{\mathbf{O}}{x},\mathbf{O}>0$ of $\mathcal{L}_{cauchy}$ is convex. Assuming a constant $\mathbf{O}\in\mathbb{R}^{+}$ and solving for $\min_{x}\Phi(x)$, we have $x^{*}=\sqrt{\mathbf{O}}$. Similarly, in Stage 2, if $\mathbf{\Gamma}^{raw}$ is randomly initialized to values largely different from $\mathbb{E}^{ACP}_{stage1}$ (the ACP error from Stage 1), convergence will be highly unstable. Therefore, we initialize $\mathbf{\Gamma}^{raw}=\mathbb{E}^{ACP}_{stage1}$, and optimize $\mathbf{P}^{cali}$, $f$, $\mathbf{Q}$, $\mathbf{t}$, and $\mathbf{\Gamma}^{raw}$ jointly.
4 Experiments
To demonstrate the superiority of our method, we show extensive quantitative and qualitative results in this section. For NeRF-DS nerfds , DAVIS davis , and iPhone iphone datasets without GT camera parameters, we feed the camera parameters from different methods to 4DGS 4dgs , while keeping all other factors the same, and evaluate each NVS performance (PSNR, SSIM, and LPIPS). Regarding the MPI-Sintel mpi-sintel and TUM-dynamics tum-dynamics datasets with GT camera parameters, we directly evaluate methods by ATE, RPE trans, and RPE rot metrics. In all tables, the best and second-best results are bold and underline. More about datasets, and evaluation metrics are in appendix˜C and appendix˜D.
[…]
Method NeRF-DS DAVIS
RoDynRFrodynrf 23.033/0.749/0.385 -
$\texttt{COL}^{\text{w/ mask}}$ 32.174/0.923/0.147 -
$\texttt{COL}^{\text{w/o mask}}$ 29.348/0.875/0.224 9.196/0.236/0.435
casualSAMcasualsam 21.230/0.686/0.463 19.032/0.486/0.482
Ours 33.552/0.938/0.118 22.292/0.709/0.279
Table 3: Runtime Evaluation on NeRF-DS nerfds , DAVIS davis , and iPhone iphone , covering frame count from 50 to 900. is supervised by additional GT priors. Our method is the most efficient.
Method NeRF-DS DAVIS iPhone
RoDynRF rodynrf 29.6h 27.4h 28.5h
$\texttt{COL}^{\text{w/ mask}}$ 1.5h - -
$\texttt{COL}^{\text{w/o mask}}$ 1.8h 0.51h 9.53h
casualSAM casualsam 10.5h 0.28h 4.07h
Ours 0.83h 0.03h 0.33h
Table 4: Camera Pose Evaluation on TUM-dynamics tum-dynamics . Other results are from Cut3r cut3r and Monst3r monst3r . Performance of DROID-SLAM droidslam is from casualSAM casualsam . Our method achieves the best overall performance among all RGB-only supervised methods, and even better than the ones supervised by additional GT priors.
[…]
GT 3D Point Cloud
& Camera Pose
Monst3r monst3r 0.098 0.019 0.935
Dust3r dust3r 0.083 0.017 3.567
Mast3r mast3r 0.038 0.012 0.448
Cut3r cut3r 0.046 0.015 0.473
GT Focal Length
   + GT Motion Prior
LEAP-VO leapvo 0.046 0.027 0.385
GT Focal Length
   + Metric Depth
DROID-SLAM droidslam 0.043 - -
GT Motion Priors ParticleSfM particlesfm - - -
RGB-Only Robust-CVD robustcvd 0.153 0.026 3.528
casualSAM casualsam 0.071 0.010 1.712
Ours 0.065 0.010 0.987
4.1 Implementation Details
The optimization is conducted on 1 NVIDIA A100 40GB GPU with Adam adam optimizer and learning rates $l_{\mathbf{Q}}=0.01$, $l_{\mathbf{t}}=0.01$, $l_{f}=1.0$, $l_{\mathbf{P}^{cali}}=0.01$, and $l_{\mathbf{\Gamma}^{raw}}=0.01$. We also choose to build our patch-wise tracking filters on CoTracker cotracker and load its pre-training weights. The hyperparameters of our patch-wise tracking filters are set at $\tau_{var}=0.1$, $B=100$, $w_{\text{NeRF-DS, DAVIS, MPI-Sintel}}=12$, and $w_{\text{iPhone, TUM}}=24$. Notably, $w$ is only related to the frame size. Besides, throughout our experiments, we have 200 and 50 iterations in $\text{Stage}_{1}$ and $\text{Stage}_{2}$ respectively.
[…]
Method alley_1 alley_2 ambush_4 ambush_5 ambush_6 market_2 market_6
casualSAM casualsam 0.028/0.006/0.057 0.003/0.003/0.392 0.040/0.058/0.321 0.053/0.040/0.211 0.302/0.088/2.362 0.010/0.010/0.041 0.239/0.207/0.544
Ours 0.002/0.003/0.038 0.009/0.002/0.047 0.119/0.049/1.367 0.065/0.039/1.192 0.080/0.129/2.191 0.003/0.010/0.110 0.009/0.006/0.301
Method shaman_3 sleeping_1 sleeping_2 temple_2 mountain_1 bamboo_1 bamboo_2
casualSAM casualsam 0.008/0.009/0.050 0.017/0.016/0.173 0.013/0.025/0.170 0.005/0.004/0.380 0.003/0.004/0.182 0.033/0.009/0.056 0.005/0.003/0.035
Ours 0.003/0.001/0.085 0.008/0.001/0.074 0.002/0.001/0.034 0.017/0.003/0.142 0.007/0.004/0.060 0.003/0.003/0.033 0.004/0.003/0.033
4.2 Time Efficiency Evaluation
[…]
Figure 5: Qualitative NVS Results on DAVIS davis . Our performance is the best because of our accurate camera estimates. More are in section˜E.2.2 (fig.˜13, fig.˜14, fig.˜15, and fig.˜16).
Figure 6: Qualitative Results of Camera Pose on MPI-Sintel mpi-sintel . represents our camera estimates; represents the GT. Our estimated camera trajectories almost perfectly align with the GT.
Besides, in Figure˜4, we see that our method exhibits a linear growth (at the rate of about 1/800 hours per frame) vs $\texttt{COL}^{w/omask}$ whose runtime growth is roughly exponential. This difference will be increasingly significant as the video length increases, which can also demonstrate the superior time efficiency of our method compared with other RGB-only supervised methods. We exclude the casualSAM here since its runtime is too large to fit here.
4.3 Camera Pose Evaluation
We follow the same evaluation setup of Cut3r cut3r and Monst3r monst3r on TUM-dynamics tum-dynamics and evaluate all videos of the synthetic MPI-Sintel mpi-sintel dataset.
[…]
Method Apple Paper-..mill Space-out Backpack Block Creeper Teddy
Record3D 26.35/0.77/0.33 23.91/0.73/0.24 27.12/0.77/0.33 20.79/0.56/0.40 23.72/0.71/0.38 21.80/0.63/0.27 19.72/0.59/0.41
$\texttt{COL}^{\text{w/o mask}}$ 22.45/0.69/0.41 22.74/0.72/0.28 24.33/0.74/0.38 18.58/0.39/0.54 18.49/0.60/0.49 18.13/0.43/0.48 16.56/0.50/0.50
casualSAMcasualsam 19.03/0.58/0.57 18.85/0.39/0.54 22.09/0.67/0.47 18.41/0.36/0.55 19.10/0.59/0.52 16.40/0.29/0.62 15.69/0.42/0.58
Ours 25.96/0.74/0.37 24.09/0.74/0.22 28.42/0.79/0.31 21.22/0.64/0.32 23.28/0.69/0.38 21.67/0.63/0.28 20.78/0.60/0.41
Method Handwavy Haru-sit Mochi-..five Spin Sriracha Pillow Wheel
Record3D 27.80/0.86/0.24 29.86/0.87/0.22 34.34/0.91/0.24 24.85/0.69/0.38 31.15/0.87/0.25 20.86/0.63/0.41 20.78/0.60/0.41
$\texttt{COL}^{\text{w/o mask}}$ 15.69/0.61/0.55 25.58/0.80/0.30 22.47/0.77/0.38 19.27/0.55/0.49 28.41/0.86/0.28 14.75/0.46/0.57 20.78/0.60/0.41
casualSAMcasualsam 20.87/0.68/0.46 19.88/0.69/0.41 26.34/0.84/0.35 19.33/0.45/0.57 23.20/0.73/0.42 16.95/0.51/0.55 14.69/0.47/0.57
Ours 28.02/0.86/0.22 28.31/0.85/0.24 34.56/0.92/0.22 24.81/0.67/0.39 32.49/0.89/0.25 20.63/0.61/0.44 20.42/0.67/0.37
Quantitative Evaluation. In Table˜5 and Section˜4, our method has the best performance among all RGB-only supervised approaches. Our method also achieves comparable or even better results than others that require additional GT priors as supervision. We attribute it to our accurate and robust pseudo-supervision, derived from RGB-only input, enabling effective outlier-aware joint optimization. Besides, our uncertainty modeling and loss design effectively down-weight the impact of moving outliers. Since $\texttt{COL}^{\text{w/ mask}}$ and $\texttt{COL}^{\text{w/o mask}}$ always fail on MPI-Sintel mpi-sintel , as observed by us and rodynrf ; monst3r , we exclude comparisons with them here. However, RGB-only supervised methods including ours perform not very well in some special cases, which is discussed in the limitations.
Qualitative Evaluation. In Figure˜6, we show our estimated camera trajectories alongside the GT on MPI-Sintel mpi-sintel . Our estimated camera trajectories can perfectly overlap with the GT, which provides qualitative support to the higher accuracies seen in the quantitative results in Table˜5.
4.4 NVS Evaluation
Since NeRF-DS nerfds , DAVIS davis , and iPhone iphone datasets do not provide GT camera parameters, we follow cf3dgs ; rodynrf ; gflow ; sc-4dgs by inputting camera estimates of different methods into the same 4D reconstruction pipeline - 4DGS 4dgs , and evaluate the NVS performance. Such NVS performance reveals the quality of the camera parameter estimation.
Figure 7: Qualitative NVS Results on iPhone iphone . Our method outperforms other SOTA RGB-only supervised approaches and even surpasses LiDAR-based Record3D when the movement in scenes with large motion (top row). More are in section˜E.2.2 (fig.˜10, fig.˜11, and fig.˜12).
Quantitative Evaluation. In LABEL:tab:nerf-ds&davis, our method is the best on NeRF-DS nerfds (long videos w/ little blur, textureless regions, and specular moving objects) and DAVIS davis (short videos w/ low parallax and rapid object movement), demonstrating our more accurate camera estimates. We skip $\texttt{COL}^{\text{w/ mask}}$ and RoDynRF on DAVIS davis because they are not RGB-only supervised methods and require supervision beyond RGB frames, and have already underperformed compared to ours on NeRF-DS nerfds . Besides, our pseudo-supervision extraction built on the PT models cotracker ; cotracker3 performs better on low-parallax videos, which remains challenging for the pre-trained depth model midas .
Regarding the iPhone iphone dataset (videos w/ irregular camera movement and object movement), it provides so-called ’GT’ camera parameters obtained by Record3D which is a paid mobile app obtaining camera results by LiDAR sensors. However, we observe that such so-called ’GT’ camera parameters are occasionally unreliable. As shown in Table˜6 and fig.˜7, besides being the best among all RGB-only supervised methods, our method can occasionally beat Record3D.
Figure 8: Qualitative NVS results on NeRF-DS nerfds . Our renderings are the most plausible. More are in section˜E.2.2 (fig.˜17).
Table 7: Ablation Study on NeRF-DS nerfds - Part 1. We conduct ablation studies on different scene optimization strategies and individual components of our proposed model.
Scene
Optimization
Camera
Optimization
[…]
D-3DGS d3dgs $\texttt{COL}^{\text{w/o mask}}$ 31.14 0.9192 0.1609
Ours (full) 32.45 0.9336 0.1312
4DGS 4dgs $\texttt{COL}^{\text{w/o mask}}$ 29.35 0.8748 0.2240
Ours (full) 33.55 0.9381 0.1182
+ w/o two-stage 25.95 0.8100 0.2668
+ w/o $\Gamma$ 26.44 0.8667 0.2327
+ w/o $\mathbb{E}^{ACP}$ 23.56 0.7203 0.3139
+ w/o texture filt. 25.99 0.8356 0.2536
+ w/o gradient filt. 26.04 0.8393 0.2404
+ w/o distrib. filt. 26.02 0.8382 0.2497
Qualitative Evaluation. We evaluate the quality of the rendered RGB images and depth maps in Figure˜7, Figure˜8, and Figure˜6. Beyond superior RGB renderings, our camera estimates yield the highest-quality depth maps, offering more convincing evidence of accurate scene geometry than RGB renderings. It indicates that our estimated camera parameters enable the model to learn the correct dynamic scene representations rather than overfitting to training views. In the first row of Figure˜7, ours performs the best (surpassing even Record3D), especially in rendered depth; whereas in the second row, our method does not match Record3D, but is still better than other RGB-only supervised works. This is because Record3D is not originally designed for dynamic scenes, so when a scene contains larger irregular movements, its performance will be worse (such observations are also supported by the numerical results in Table˜6). In contrast, our method is more robust in various scenarios, consistently maintaining high standards.
Table 8: Ablation Study on NeRF-DS nerfds - Part 2. We conducted ablation studies on different PT models.
[…]
4DGS 4dgs Ours + built on CoTracker cotracker 33.55 0.9381 0.1182
Ours + built on CoTracker3 cotracker3 33.52 0.9384 0.1180
Ablation Study. In Figure˜8, the loss of any filter results in less robust relations across video, leading to poor camera estimates and NVS performance. Further, the removal of any of $\mathbf{\Gamma}$, $\mathbb{E}^{ACP}$, or the two-stage strategy will harm the results due to outliers. This indicates that w/o such a strategy, increasing training iterations is a necessary but not sufficient condition for comparable results. We also overcome the limitations of COLMAP colmap by improving the performance of different scene optimization models d3dgs ; 4dgs with our camera estimates. As reported in CoTracker3 cotracker3 , CoTracker cotracker performs worse than CoTracker3. However, in table˜8, the performance of our proposed method is nearly independent of building on the particular PT model. This further supports our claim that our patch-wise tracking filters effectively exact only the accurate trajectories as pseudo-supervision.
5 Conclusion and Limitation
We proposed a new RGB-only supervised, accurate, and efficient camera parameter optimization method in casually collected dynamic-scene videos. Our method effectively tackles the challenge of precisely and efficiently estimating per-frame camera parameters under the situation of having no additional GT supervision (e.g. GT motion masks, focal length, 3D point clouds, metric depth, and camera poses) other than RGB videos, which is the most common scenario in real-world and consumer-grade reconstructions. Our method may serve as a step towards high-fidelity dynamic scene reconstruction from casually captured videos. Due to the space limit, the limitation is discussed in appendix˜F.
Appendix A Dynamic Scene Optimization
Preliminary. As an alternative to NeRF nerf , which has a lengthy runtime, 3DGS 3dgs recently introduced a new way to learn static scene representations in terms of explicit 3D Gaussian ellipsoids. Unlike the implicit representations of NeRF stored as the weights in the Convolutional Neural Network (CNN), 3DGS 3dgs uses explicit representations in 3D world coordinates and performs differential Gaussian rasterization on GPUs using CUDA which significantly speeds up computational efficiency. Each 3D Gaussian ellipsoid $\mathcal{G}$ (x) is parameterized by its (1) Gaussian center $\mathcal{X}\in\mathbb{R}^{3}$; (2) quaternion factor $\mathbf{r}\in\mathbb{R}^{4}$; (3) opacity $\alpha\in\mathbb{R}$; (4) scaling factor $\mathbf{s}\in\mathbb{R}^{3}$; and (5) color $\mathcal{C}\in\mathbb{R}^{k}$ (k denotes degrees of freedom), and represented by:
$\mathcal{G}(x)=e^{-1/2(x-\mu)^{T}\Sigma^{-1}(x-\mu)}\\$ (9)
$\Sigma^{\prime}=\mathbf{JW}\Sigma\mathbf{W}^{T}\mathbf{J}^{T},\hskip 6.0pt\Sigma=\mathbf{\mathbf{r}\mathbf{s}}\mathbf{s}^{T}\mathbf{r}^{T}$ (10)
where $\Sigma$ is the 3D covariance matrix in the world space, $\mathbf{W}$ and $\mathbf{J}$ are the view transformation matrix and the Jacobian matrix of the affine transformation parts, respectively, of the projective transformation, and $\Sigma^{\prime}$ is the covariance matrix in the camera coordinates. The color is rendered by:
$\mathcal{C}(p)=\sum_{k\in K}c_{k}\alpha_{k}\Pi_{k}^{j-1}(1-\alpha_{k})$ (11)
where $c_{k}$ and $\alpha_{k}$ represent the spherical harmonic (SH) coefficient and the density at this point.
4D GS. To take advantage of its optimization efficiency, we use 4DGS 4dgs to learn dynamic scene representations. With different camera estimates and the same scene optimization method 4dgs , we use NVS performance to evaluate the accuracy of camera parameter estimates. In 4DGS4dgs , the NVS performance depends on how well the canonical representations and deformation representations are optimized. The canonical (refers to ’mean’ as in the previous work 3dgs ; d3dgs ; 4dgs ) representations $\mathcal{G}$ are learned by a canonical Gaussian field to optimize the mean (canonical) position $\mathcal{X}\in\mathbb{R}^{3}$, color $\mathcal{C}\in\mathbb{R}^{k}$, opacity $\alpha\in\mathbb{R}$, quaternion factor $r\in\mathbb{R}^{4}$, scaling factor $s\in\mathbb{R}^{3}$, and the deformation representations $\mathcal{F}$ are optimized using a deformation field 4dgs to learn the offsets $\Delta\mathcal{G}$, supervised by an L1 loss between images and renderings. Since the color and opacity of the Gaussian ellipsoids do not change over time, the deformed attributes consist of $(\mathcal{X}^{\prime},r^{\prime},s^{\prime})=(\mathcal{X}+\Delta\mathcal{X},r+\Delta r,s+\Delta s)$. More details can be found in 4DGS 4dgs .
Appendix B Derivation of Cauchy Negative-log-likelihood
The Cauchy loss function is derived from Equation˜12 to Equation˜14. Since we use the Cauchy probability density function (PDF) to model the uncertainty of calibration points $\mathbf{P^{cali}}$, we want to maximize the likelihood of the Cauchy PDF:
$f(x;x_{0},\mathbf{\Gamma})=\frac{1}{\pi\mathbf{\Gamma}[1+(\frac{x-x_{0}}{\mathbf{\Gamma}})^{2}]},\quad\mathbf{\Gamma}>0$ (12)
Equivalently, we minimize the negative-log-likelihood of $f(x;x_{0},\mathbf{\Gamma})$, to define the loss function::
$\text{NLL}(x;x_{0},\mathbf{\Gamma})$ $=-\log(f(x;x_{0},\mathbf{\Gamma}))$ (13)
$=\log(\pi\mathbf{\Gamma})+\log(1+(\frac{x-x_{0}}{\mathbf{\Gamma}})^{2})$
$=\log[\pi\cdot(\mathbf{\Gamma}+\frac{(x-x_{0})^{2}}{\mathbf{\Gamma}})]$
$=\log\pi+\log(\mathbf{\Gamma}+\frac{(x-x_{0})^{2}}{\mathbf{\Gamma}})$
where $\log\pi$ and $x_{0}$ denote a constant term and the ground truth which can be omitted. Thus, our objective is as follows:
$\min_{x,\mathbf{\Gamma}}\text{NLL}(x;x_{0},\mathbf{\Gamma})=\min_{x,\mathbf{\Gamma}}\log(\mathbf{\Gamma}+\frac{(x-x_{0})^{2}}{\mathbf{\Gamma}})$ (14)
Appendix C Datasets
To demonstrate our performance on a broader range of scenarios, we have conducted extensive experiments across five public datasets - NeRF-DS nerfds , DAVIS davis , iPhone iphone , MPI-Sintel mpi-sintel , and TUM-dynamics tum-dynamics . These videos contain different camera and object motion patterns, and different texture levels. The lengths of the videos range from about 50 to 900. Regarding the train/test split of the NVS evaluation, for every 2 adjacent frames, we take the first frame for training and the second frame for testing. For the setup of camera pose evaluation, we follow Cut3r cut3r and Monst3r monst3r in the experiments on TUM-dynamics and evaluate all videos in MPI-Sintel mpi-sintel .
NeRF-DS. NeRF-DS nerfds dataset includes seven long monocular videos (400-800 frames) of different dynamic, real-world indoor scenarios with little blur. Each video has at least one specular moving object against a mix of low-texture and high-texture backgrounds. NeRF-DS nerfds exhibits large scene and camera movements, so the frames have some blur. The GT motion masks provided are human-labeled and the camera parameters are estimated by $\texttt{COL}^{\text{w/ mask}}$. Like previous works nerfds ; d3dgs , we take the highest resolution images available (480 $\times$ 270) as the RGB input in all experiments.
DAVIS. DAVIS davis dataset contains 40 short monocular videos that capture different dynamic scenes in the wild. Each video has 50-100 frames, including at least one dynamic object. The GT motion masks are also provided as in NeRF-DS nerfds . However, like  rodynrf ; gflow ; casualsam ; particlesfm , we exclude some videos using fixed cameras, changeable focal lengths, etc. Different from others rodynrf ; gflow ; casualsam ; leapvo which only show experiments of about 10 videos in DAVIS davis , we conduct experiments on 21 videos containing large camera and object movements. We utilize the RGB frames with the resolution of 854 $\times$ 480 as input.
iPhone. The iPhone iphone dataset is an extremely challenging dataset (180-475 frames) with significant camera rotations and translations, and rapid movements of objects. There are 14 monocular videos including indoor and outdoor scenes and no GT motion mask is provided. It would also be difficult to insert motion masks for this dataset because there is no clear boundary between the moving and stable regions within any frame. They represent real-world casually recorded videos. We conduct experiments on all of them. The frame size is 720 $\times$ 960. These videos are recorded by the Record3D record3d app on iPhone which uses LiDAR sensors to obtain metric depth for camera estimation. In our comparisons with the camera estimates provided by Record3D, we also take the Record3D app as one of the baselines and compare with it.
MPI-Sintel. MPI-Sintel mpi-sintel is a synthetic dataset provided GT camera parameters. It has 18 short videos (about 50 frames) in total containing large object movement. In some cases, the moving objects cover most of the screen. Most of the existing works rodynrf ; monst3r ; leapvo select 14 videos for evaluation, but in this paper, we evaluate the methods among all the videos. The synthetic MPI-Sintel dataset exhibits domain gaps compared to real-world scenarios, which is considered to be one of the reasons why some existing methods casualsam ; particlesfm perform well on MPI-Sintel, but do not work efficiently on other real-world datasets. In experiments, we take the frames with default sizes as the input the our method, while keeping the default resizing setup of the other methods.
TUM-dynamics. TUM-dynamics tum-dynamics dataset contains 8 long real-world blurry videos recording the dynamic indoor scenes provided with GT camera parameters. However, although the videos in this dataset are indoor scenes, each video features a significant depth of field. TUM-dynamics dataset also contains large camera movement and rapid object movement. We follow the experimental setup of MonST3R monst3r on this dataset, which is sampling the first 90 frames with the temporal stride of 3 to save compute.
Appendix D Evaluation Metrics
As discussed in section˜4 of the main paper, we directly conduct camera pose evaluation against the GT on MPI-Sintelmpi-sintel and TUM-dynamics tum-dynamics , using the standard metrics: ATE, RPE trans, and RPE rot. Besides, regarding NeRF-DS nerfds , DAVIS davis , and iPhone iphone datasets which are not provided with GT camera parameters, we conduct NVS evaluation with standard metrics: PSNR, SSIM, and LPIPS. We also employ time evaluation to demonstrate the superior time efficiency of our method.
D.1 PSNR & SSIM & LPIPS
PSNR. PSNR is a measure of the ratio between the maximum possible power of a signal and the power of corrupting noise that affects the fidelity of its representation. PSNR is commonly used to compare the qualities of the original and the rendered images, and is obtained from the Mean Square Error (MSE) between the original and the rendered images:
$\text{PNSR}=10\cdot\log_{10}\left(\frac{\text{MAX}^{2}}{\text{MSE}(\text{Image}_{\text{Rendered}},\text{Image}_{\text{GT}})}\right)$ (15)
, where MAX is the maximum pixel value of the image.
SSIM. SSIM measures the similarity between two images based on structural information. Its evaluation involves luminance, contrast, and structure. Compared to PSNR, SSIM is intended to match human perception more closely. The SSIM values range from -1 to 1, where 1 denotes perfect. It is given as:
$\text{SSIM}=\frac{(2\mu_{x}\mu_{y}+c_{1})(2\sigma_{xy}+c_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+c_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+c_{2})}$ (16)
where $x$ and $y$ are two images, $\mu_{x}$, $\mu_{y}$ and $\sigma_{x}^{2}$, $\sigma_{y}^{2}$ are the corresponding averages and variances of $x$ and $y$, $\sigma_{xy}$ represents the covariance of $x$ and $y$, and $c_{1}$ and $c_{2}$ denote the regularization terms.
LPIPS. LPIPS measures perceptual similarity in terms of features of deep neural networks, such as pre-trained VGG vgg or AlexNet alexnet . It compares feature activations of image patches. Like 3dgs ; d3dgs ; 4dgs ; sc-4dgs , we here use the VGG-based networks.
D.2 ATE & RPE trans & RPE rot
ATE. ATE quantifies the difference between the actual trajectory and the estimated trajectory of a robot or camera over time, offering a global measure of error along the entire path. It is calculated by aligning the estimated trajectory with the ground truth and then measuring the Euclidean distance between each corresponding point on the two trajectories.
RPE trans. RPE trans quantifies the error in the translational component between consecutive poses or over a fixed time/distance interval. Unlike ATE, which assesses the overall trajectory, RPE Trans emphasizes the local accuracy of the motion estimation by evaluating how well the system preserves the relative motion between two points in time or space.
RPE rot. RPE rot quantifies the error in the orientation component between the estimated poses and the ground truth. This metric is computed by measuring the difference in orientation over short sequences, and it is typically expressed in angular units, such as degrees or radians.
Appendix E More Results
E.1 Quantitative Results
E.1.1 Runtime
We report the detailed runtime comparisons on the NeRF-DS nerfds , DAVIS davis , and iPhone iphone datasets, each containing over 50 frames, where runtime differences become more pronounced. Specifically, in Table˜9 and Table˜10, the runtime of our method is the shortest. In addition, in Table˜11, on the NeRF-DS nerfds dataset, the runtime of $\texttt{COL}^{\text{w/ mask}}$ and $\texttt{COL}^{\text{w/o mask}}$ on Plate video is shorter than of ours. This is because $\texttt{COL}^{\text{w/ mask}}$ and $\texttt{COL}^{\text{w/o mask}}$ fail on this video, leading to a quick convergence to the local minima. Such a conclusion can also supported by qualitative results of fig.˜8 in the main paper.
Table 9: Quantitative Runtime Results on DAVIS davis . Cam $\rightarrow$ camera optimization time; Cam+Scene $\rightarrow$ overall (camera+scene) optimization time; h $\rightarrow$ hour; m $\rightarrow$ minute. We mark the shortest time in bold. Our method is the most efficient without any failure.
Method Ours $\texttt{COL}^{\text{w/o mask}}$ casualSAM
Cam Cam+Scene Cam Cam+Scene Cam Cam+Scene
Camel 1.57m 25.28m 40m 71m 24m 46m
Bear 3.15m 1.08h 56m 88m 20m 39m
Breakdance-flare 1.73m 0.92h FAIL - 15m 34m
Car-roundabout 4.97m 21.95m 10m 41m 18m 46m
Car-shadow 0.93m 17.88m 5m 31m 10m 36m
Car-turn 2.97m 17.35m FAIL - 27m 49m
Cows 2.85m 22.50m 73m 84m 26m 48m
Dog 1.60m 11.70m 10m 53m 12m 31m
Dog-agility 0.67m 26.63m FAIL - 5m 31m
Goat 2.10m 18.95m 107m 124m 22m 44m
Hike 4.20m 31.83m FAIL - 20m 42m
Horsejump-high 1.97m 21.20m 5m 33m 10m 31m
Lucia 1.97m 22.75m 44m 65m 16m 36m
Motorbike 2.08m 18.77m 6m 31m 9m 32m
Parkour 9.07m 26.65m 16m 37m 27m 48m
Rollerblade 1.25m 17.08m FAIL - 8m 27m
Tennis 3.47m 17.03m 9m 30m 17m 38m
Train 1.90m 18.22m 32m 57m 19m 44m
Mean 2.68m 21.02m 31m 56m 17m 39m
Table 10: Quantitative Runtime Results on iPhone iphone . Cam $\rightarrow$ camera optimization time; Cam+Scene $\rightarrow$ overall (camera+scene) optimization time; h $\rightarrow$ hour; m $\rightarrow$ minute. We mark the shortest time in bold. Our method is the most efficient.
Method Ours $\texttt{COL}^{\text{w/o mask}}$ casualSAM
Cam Cam+Scene Cam Cam+Scene Cam Cam+Scene
Apple 33m 47m 10.95h 11.25h 6.80h 7.32h
Paper-windmill 15m 27m 8.18h 8.70h 3.50h 3.97h
Space-out 23m 69m 4.42h 4.72h 5.87h 6.30h
Backpack 7m 25m 1.83h 2.12h 1.58h 2h
Block 27m 42m 10h 10.45h 6h 6.50h
Creeper 27m 45m 16.03h 16.45h 4.38h 4.82h
Handwavy 15m 30m 4.62h 5.1h 3.30h 3.72h
Haru-sit 10m 25m 0.77h 1.18h 1.25h 1.72h
Mochi-high-five 6m 18m 0.67h 1.03h 1.53h 1.93h
Pillow 21m 36m 19.18h 19.70h 3.70h 4.20h
Spin 30m 44m 20h 20.58h 5.50h 6h
Sriracha-tree 15m 27m 4.58h 4.88h 3.22h 3.68h
Teddy 31m 46m 19h 19.71h 6.78h 7.28h
Wheel 27m 46m 6m 13.80h 3.67h 4.17h
Mean 20m 38m 9.53h 9.97h 4.07h 4.53h
Table 11: Quantitative runtime results on NeRF-DS nerfds . Cam $\rightarrow$ camera optimization time; Cam+Scene $\rightarrow$ overall (camera+scene) optimization time; h $\rightarrow$ hour; m $\rightarrow$ minute. We mark the shortest time in bold. we show only Cam+Scene of RoDynRF rodynrf due to its joint optimization of the camera and scene. is supervised by additional GT priors. $\texttt{COL}^{w/mask}$ and $\texttt{COL}^{w/omask}$ are faster than us on Plate because they fail on this video, leading to a quick convergence to the local minima, which can be supported by qualitative results of fig.˜8 in the main paper. Among all, our method is the most efficient.
Method Ours $\texttt{COL}^{\text{w/ mask}}$ $\texttt{COL}^{\text{w/o mask}}$ casualSAM RoDynRF
Cam Cam+Scene Cam Cam+Scene Cam Cam+Scene Cam Cam+Scene Cam+Scene
Bell 1.05h 1.20h 2.50h 2.72h 3.00h 3.25h 16.5h 16.8h 28.6h
As 0.95h 1.08h 2.00h 2.17h 2.55h 2.72h 14.87h 15.08h 33.6h
Basin 0.75h 0.92h 1.42h 1.62h 1.60h 1.85h 9.88h 10.67h 33.8h
Plate 0.53h 0.68h 0.42h 0.60h 0.50h 0.87h 4.67h 4.98h 25.6h
Press 0.68h 0.82h 0.85h 1.05h 0.90h 1.08h 6.28h 6.60h 28.5h
Cup 1.02h 1.15h 2.37h 2.58h 2.57h 2.73h 13.50h 13.78h 28.8h
Sieve 0.78h 0.92h 1.15h 1.35h 1.58h 1.77h 7.83h 8.13h 28.3h
Mean 0.83h 0.97h 1.52h 1.73h 1.82h 2.03h 10.50h 10.80h 29.6h
E.2 Qualitative Results
E.2.1 Trajectories from Patch-wise Tracking Filters as Pseudo-supervision
In Figure˜9, we show the trajectory comparisons on the NeRF-DS nerfds dataset as samples. As discussed in the 3rd paragraph in section˜1, and section˜3.1 of the main paper, our patch-wise tracking filters can establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision, avoiding noisy and inaccurate tracking trajectories. In Figure˜9, it is easy to see our proposed method avoids the inaccurate ones in the low-texture regions (walls), and meanwhile, adaptively adds new reliable trajectories when the number of left trajectories on each frame is less than $B$.
Figure 9: Trajectory Comparisons on the NeRF-DS nerfds Dataset. In each scenario, top row $\rightarrow$ $F_{0}$; bottom row $\rightarrow$ $F_{247}$; $w/o$ patch-wise filters $\rightarrow$ raw CoTracker cotracker ; $w/$ patch-wise filters $\rightarrow$ Ours. It is easy to see our proposed method avoids the inaccurate trajectories in the low-texture regions, whereas the trajectories of the points in the low-texture regions tracked by raw CoTracker cotracker are extremely unreliable.
E.2.2 NVS
We show more RGB and depth rendering results on NeRF-DS nerfds , DAVIS davis , and iPhone iphone dataset in Figure˜10, Figure˜11, Figure˜12, Figure˜13, Figure˜14, Figure˜15, Figure˜16, and Figure˜17. It is easy to see that the RGB and depth rendering results of our method are better than other RGB-only supervised approaches. In addition, the performance of our method is also comparable with that of the LiDAR-based Record3D app.
Figure 10: More Qualitative NVS Results on iPhone iphone - Part 1. Our renderings exhibit higher fidelity and more accurate geometry compared to other RGB-only supervised methods. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Figure 11: More Qualitative NVS Results on iPhone iphone - Part 2. Our renderings exhibit higher fidelity and more accurate geometry compared to other RGB-only supervised methods. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Figure 12: More Qualitative NVS Results on iPhone iphone - Part 3. Our renderings exhibit higher fidelity and more accurate geometry compared to other RGB-only supervised methods. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Figure 13: More Qualitative NVS Results on DAVIS davis - Part 1. Our renderings exhibit higher fidelity compared to other RGB-only supervised methods.
Figure 14: More Qualitative NVS Results on DAVIS davis - Part 2. Our renderings exhibit higher fidelity compared to other RGB-only supervised methods.
Figure 15: More Qualitative NVS Results on DAVIS davis - Part 3. Our renderings exhibit higher fidelity compared to other RGB-only supervised methods.
Figure 16: More Qualitative NVS Results on DAVIS davis - Part 4. Our renderings exhibit higher fidelity compared to other RGB-only supervised methods.
Figure 17: More Qualitative NVS Results on NeRF-DS nerfds . Our renderings exhibit higher fidelity compared to other RGB-only supervised methods.
E.2.3 Optimized 3D Gaussian Fields
Since the iPhone iphone dataset is the most challenging dataset with large camera and object movements, we show more visualizations of optimized 3D Gaussian fields in Figure˜19, Figure˜20, and Figure˜18. Such comparisons demonstrate that our camera estimates enable superior reconstruction of 3D Gaussian fields compared to other RGB-only supervised approaches. Moreover, the reconstructed fields using our estimates are comparable to, or even surpass, those obtained with the LiDAR-based Record3D record3d app, particularly in scenes with significant motion.
Figure 18: Optimized 3D Gaussian Fields on iPhone iphone - Part 1. Our reconstructed 3D Gaussian Fields are more geometrically accurate compared to the ones of other RGB-only supervised methods, which demonstrates our camera estimates are more accurate. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Figure 19: Optimized 3D Gaussian Fields on iPhone iphone - Part 2. Our reconstructed 3D Gaussian Fields are more geometrically accurate compared to the ones of other RGB-only supervised methods, which demonstrates our camera estimates are more accurate. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Figure 20: Optimized 3D Gaussian Fields on iPhone iphone - Part 3. Our reconstructed 3D Gaussian Fields are more geometrically accurate compared to the ones of other RGB-only supervised methods, which demonstrates our camera estimates are more accurate. Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
Appendix F Limitation
Although our proposed method is currently the most state-of-the-art RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes, there are still several limitations. We assume a constant focal length throughout the video. While this assumption is reasonable and currently common to SOTA, the task of accurate and efficient camera parameter optimization for dynamic scene videos with zooming effects under RGB-only supervision remains an open problem. Another common challenge for RGB-only supervised methods, not addressed in this paper, is maintaining robustness in scenes dominated by large moving objects. As shown in fig.˜21, the screen space is occupied by the moving human and dragon. It is challenging for our method to establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision because most of the extracted trajectories belong to outliers. CasualSAM casualsam struggles due to the rapid changes in depth maps from frame to frame, making 3D space alignment difficult. We plan to maintain consistency in our input setup and address these challenges as part of future research.
Figure 21: Failure cases of ours and casualSAM on MPI-Sintel mpi-sintel .

### v2

@@
 Many efforts attempted to improve it by incorporating more priors as supervision such as GT focal length, motion masks, 3D point clouds, camera poses, and metric depth, which, however, are typically unavailable in casually captured RGB videos.
-In this paper, we propose a novel method for more accurate and efficient camera parameter optimization in dynamic scenes solely supervised by a single RGB video.
+In this paper, we propose a novel method for more accurate and efficient camera parameter optimization in dynamic scenes solely supervised by a single RGB video, dubbed ROS-Cam.
 Our method consists of three key components: (1) Patch-wise Tracking Filters, to establish robust and maximally sparse hinge-like relations across the RGB video.
@@
 […]
-Based on these insights, we propose an RGB-only supervised, accurate, and efficient camera parameter optimization method, with a brief performance overview in Figure˜1.
+Based on these insights, we propose ROS-Cam, an RGB-only supervised, accurate, and efficient camera parameter optimization method, with a brief performance overview in Figure˜1.
 Specifically, to minimize reliance on pre-trained dense prediction models while still establishing robust and maximally sparse hinge-like relations across the video as accurate pseudo-supervision (bottom right corner in Figure˜2), we propose the novel patch-wise tracking filters built solely on a pre-trained point tracking (PT) model.
@@
 While tracking such identified points, the visibility filter keeps removing trajectories that become invisible and the patch-wise distribution filter keeps the one with the largest gradient norm when multiple moving points enter the same patch.
-As shown in section˜E.2.1 (fig.˜9), our method only retains the robust and accurate trajectories as pseudo-supervision.
+As shown in section˜E.2.1 (fig.˜10), our method only retains the robust and accurate trajectories as pseudo-supervision.
 Figure 2: Patch-wise tracking filters.
@@
 […]
-Figure 5: Qualitative NVS Results on DAVIS davis . Our performance is the best because of our accurate camera estimates. More are in section˜E.2.2 (fig.˜13, fig.˜14, fig.˜15, and fig.˜16).
+Figure 5: Qualitative NVS Results on DAVIS davis . Our performance is the best because of our accurate camera estimates. More are in section˜E.2.2 (fig.˜14, fig.˜15, fig.˜16, and fig.˜17).
 Figure 6: Qualitative Results of Camera Pose on MPI-Sintel mpi-sintel . represents our camera estimates; represents the GT. Our estimated camera trajectories almost perfectly align with the GT.
@@
 Our method outperforms other SOTA RGB-only supervised approaches and even surpasses LiDAR-based Record3D when the movement in scenes with large motion (top row).
-More are in section˜E.2.2 (fig.˜10, fig.˜11, and fig.˜12).
+More are in section˜E.2.2 (fig.˜11, fig.˜12, and fig.˜13).
 Quantitative Evaluation.
@@
 As shown in Table˜6 and fig.˜7, besides being the best among all RGB-only supervised methods, our method can occasionally beat Record3D.
-Figure 8: Qualitative NVS results on NeRF-DS nerfds . Our renderings are the most plausible. More are in section˜E.2.2 (fig.˜17).
+Figure 8: Qualitative NVS results on NeRF-DS nerfds . Our renderings are the most plausible. More are in section˜E.2.2 (fig.˜18).
 Table 7: Ablation Study on NeRF-DS nerfds - Part 1. We conduct ablation studies on different scene optimization strategies and individual components of our proposed model.
@@
 Our method may serve as a step towards high-fidelity dynamic scene reconstruction from casually captured videos.
-Due to the space limit, the limitation is discussed in appendix˜F.
+Although our proposed method is currently the most state-of-the-art RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes, there are still several limitations.
+We assume a constant focal length throughout the video.
+While this assumption is reasonable and currently common to SOTA, the task of accurate and efficient camera parameter optimization for dynamic scene videos with zooming effects under RGB-only supervision remains an open problem.
+Another common challenge for RGB-only supervised methods, not addressed in this paper, is maintaining robustness in scenes dominated by large moving objects.
+As shown in fig.˜9, the screen space is occupied by the moving human and dragon.
+It is challenging for our method to establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision because most of the extracted trajectories belong to outliers.
+CasualSAM casualsam struggles due to the rapid changes in depth maps from frame to frame, making 3D space alignment difficult.
+We plan to maintain consistency in our input setup and address these challenges as part of future research.
+Figure 9: Failure cases of ours and casualSAM on MPI-Sintel mpi-sintel .
 Appendix A Dynamic Scene Optimization
@@
 E.2.1 Trajectories from Patch-wise Tracking Filters as Pseudo-supervision
-In Figure˜9, we show the trajectory comparisons on the NeRF-DS nerfds dataset as samples.
+In Figure˜10, we show the trajectory comparisons on the NeRF-DS nerfds dataset as samples.
 As discussed in the 3rd paragraph in section˜1, and section˜3.1 of the main paper, our patch-wise tracking filters can establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision, avoiding noisy and inaccurate tracking trajectories.
-In Figure˜9, it is easy to see our proposed method avoids the inaccurate ones in the low-texture regions (walls), and meanwhile, adaptively adds new reliable trajectories when the number of left trajectories on each frame is less than $B$.
+In Figure˜10, it is easy to see our proposed method avoids the inaccurate ones in the low-texture regions (walls), and meanwhile, adaptively adds new reliable trajectories when the number of left trajectories on each frame is less than $B$.
 Figure 10: Trajectory Comparisons on the NeRF-DS nerfds Dataset.
@@
 E.2.2 NVS
-We show more RGB and depth rendering results on NeRF-DS nerfds , DAVIS davis , and iPhone iphone dataset in Figure˜10, Figure˜11, Figure˜12, Figure˜13, Figure˜14, Figure˜15, Figure˜16, and Figure˜17.
+We show more RGB and depth rendering results on NeRF-DS nerfds , DAVIS davis , and iPhone iphone dataset in Figure˜11, Figure˜12, Figure˜13, Figure˜14, Figure˜15, Figure˜16, Figure˜17, and Figure˜18.
 It is easy to see that the RGB and depth rendering results of our method are better than other RGB-only supervised approaches.
@@
 E.2.3 Optimized 3D Gaussian Fields
-Since the iPhone iphone dataset is the most challenging dataset with large camera and object movements, we show more visualizations of optimized 3D Gaussian fields in Figure˜19, Figure˜20, and Figure˜18.
+Since the iPhone iphone dataset is the most challenging dataset with large camera and object movements, we show more visualizations of optimized 3D Gaussian fields in Figure˜20, Figure˜21, and Figure˜19.
 Such comparisons demonstrate that our camera estimates enable superior reconstruction of 3D Gaussian fields compared to other RGB-only supervised approaches.
@@
 Besides, our performance is comparable with, or even better than, the ones of the LiDAR-based Record3D app.
-Appendix F Limitation
-Although our proposed method is currently the most state-of-the-art RGB-only supervised, accurate, and efficient camera parameter optimization method in dynamic scenes, there are still several limitations.
-We assume a constant focal length throughout the video.
-While this assumption is reasonable and currently common to SOTA, the task of accurate and efficient camera parameter optimization for dynamic scene videos with zooming effects under RGB-only supervision remains an open problem.
-Another common challenge for RGB-only supervised methods, not addressed in this paper, is maintaining robustness in scenes dominated by large moving objects.
-As shown in fig.˜21, the screen space is occupied by the moving human and dragon.
-It is challenging for our method to establish robust and maximally sparse hinge-like relations as accurate pseudo-supervision because most of the extracted trajectories belong to outliers.
-CasualSAM casualsam struggles due to the rapid changes in depth maps from frame to frame, making 3D space alignment difficult.
-We plan to maintain consistency in our input setup and address these challenges as part of future research.
-Figure 21: Failure cases of ours and casualSAM on MPI-Sintel mpi-sintel .